# Reconstruction of gravitational form factors using generative machine learning

### v1

@@
We develop a generative framework based on denoising diffusion for the model-independent reconstruction of hadronic form factors from sparse and noisy data. The generative prior is built from a large ensemble of synthetic curves drawn from ten distinct functional classes rooted in different theoretical approaches to hadron structure. Applied to the proton gravitational form factors $A(t)$, $J(t)$, and $D(t)$, the framework yields non-parametric reconstructions consistent with lattice QCD across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$, remaining robust even when only one or two conditioning points are retained. The densely sampled output enables a direct extraction of the chiral low-energy constants $c_{8}=-4.6\pm 0.8$ and $c_{9}=-0.61\pm 0.19$. Using these values at the physical pion mass, we obtain $D(0)=-4.3\pm 0.8$ for the nucleon $D$-term.
@@ I Introduction
The hadronic matrix elements of the energy-momentum tensor (EMT) have been studied intensively over the past decades, both theoretically and experimentally. While much attention has been devoted to the physical interpretation of the Fourier transforms of the gravitational form factors (GFFs) in terms of spatial density distributions of energy (mass), spin, pressure, and shear forces inside hadrons, see, e.g., Jaffe:2020ebz ; Panteleeva:2022uii ; Freese:2021mzg ; Ji:2025qax , the GFFs themselves, as functions of momentum transfer, are interesting in their own right. In particular, they can be extracted from generalized parton distributions (GPDs) Muller:1994ses ; Ji:1996ek ; Radyushkin:1996nd ; Goeke:2001tz , which are experimentally accessible through exclusive processes such as deeply virtual Compton scattering (DVCS) Ji:1996ek ; Radyushkin:1997ki and hard exclusive meson production Collins:1996fb . Knowing the GFFs of a system that can be probed experimentally, such as the nucleon, establishes a direct connection between theoretical predictions and experiment, thereby testing the validity and consistency of the underlying theory.
@@
The remainder of this paper is organised as follows. In Sec. II we describe the construction of the training dataset from ten classes of physically motivated functional forms and their convex combinations. Section III presents the DDPM architecture, including the $v$-prediction formulation, the dual conditioning mechanism, and the error-propagation scheme. In Sec. IV we demonstrate the model-independent reconstruction of the proton GFFs at unphysical pion mass and extract the LECs $c_{8}$ and $c_{9}$. Details of the network architecture, training procedure, ablation studies on synthetic test curves, and the reconstruction of the GFFs at physical pion mass are collected in Appendices AD.
@@ III.3 Conditioning and error propagation
Our goal is to reconstruct the full GFF curve given $K$ known data points [^3] Here $F_{k}$ refers to a (gravitational) form factor $F(t_{k})$. $\{(t_{k},\,F_{k}\pm\sigma_{k})\}_{k=1}^{K}$ from, e.g., lattice QCD or ChPT. The model conditions on this information through two complementary mechanisms. First, we have the concatenation conditioning: each known value is mapped to the nearest grid index $j_{k}$ and normalised to the model’s internal representation: $\tilde{F}_{k}=(F_{k}-\mu_{j_{k}})/\sigma_{j_{k}}$, where $\mu$ and $\sigma$ are the per-grid-point mean and standard deviation of the training set. The network receives three input channels at every reverse step: the current noisy estimate $x_{\tau}$, a binary mask $m\in\{0,1\}^{L}$ with $m_{j_{k}}=1$ at known positions, and a condition vector $c\in\mathbb{R}^{L}$ with $c_{j_{k}}=\tilde{F}_{k}$ and zero elsewhere. During training, random masks (50% uniformly random, 30% clustered in the low-$|t|$ region, 20% unconditional) teach the network to reconstruct missing values conditioned on an arbitrary subset of known points. To propagate the uncertainties of the known data into the reconstructed uncertainty band, each of the $N$ generated samples receives its own realisation of the conditioning values,
@@ IV Model-independent reconstruction of the proton GFFs at unphysical pion mass
In this section, we demonstrate that the diffusion-based generative framework yields a fully model-independent reconstruction of the total proton GFFs $A(t)$, $J(t)$, and $D(t)$, consistent with the lattice QCD results of Ref. Hackett:2023nkr across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$, without imposing any functional form on the $t$-dependence.
The only inputs to the reconstruction are:
@@
Having learned the space of physically admissible form-factor shapes from eight distinct functional classes and their convex combinations (Sec. II), the diffusion model samples from the posterior distribution consistent with the supplied conditioning input, yielding a non-parametric estimate of both the central value and the pointwise uncertainty of the reconstructed form factor.
@@
It is worth emphasizing that for each available lattice dataset we select as conditioning input those points with only modest statistical uncertainties compared with the others, in other words, the reconstruction is not aided by large error bars that would trivially accommodate many functional shapes. The diffusion model must reproduce sharply defined values around the conditioning points while simultaneously generating physically sensible interpolations and extrapolations throughout the kinematic range. That it succeeds with as few as one or two data points is a direct consequence of the physically informed prior built into the training data, and constitutes a stringent validation of the approach.
@@ IV.1 Reconstruction of symmetry-constrained form factors $A(t)$ and $J(t)$
  • Reconstruction from minimal input: The most striking result appears in the bottom panels, where only a single lattice data point at intermediate t-t and the forward-limit constraint are supplied. Even in this extreme scenario the reconstruction yields a smoothly decaying curve whose median remains compatible with the full parametric fits. At $t=2\penalty 10000\ \mathrm{GeV}^{2}$ the credible band remains moderate for both form factors. The notably narrow uncertainty obtained with only two conditioning inputs, despite the wide range of mathematically allowed curves through those points, is a direct consequence of the physical prior encoded in the training data.
@@ IV.2 Reconstruction of D form factor
  • Minimal constraint ($D(0)<0$): The requirement that $D(0)$ be negative follows from the mechanical stability of the proton Polyakov:2018zvc ; Perevalova:2016dln : a positive $D$-term would imply a repulsive normal force and the system would collaps. This negativity is also supported by ChPT Gegelia:2021wnj , chiral quark-soliton models Goeke:2007fp ; Cebulla:2007ei , and the lattice data themselves. This weakest condition allows the form factor to fluctuate mildly at large t-t.
  • Full negativity ($D(t)<0$): Negativity is enforced at every point on the spacelike grid. This stronger condition is motivated by all known model calculation, see e.g. Refs. Cebulla:2007ei ; Cao:2025ncp ; Yao:2024ixu ; Cao:2025dkv .
@@ IV.3 Extracting the unknown LECs $c_{8}$ and $c_{9}$
The low-energy constant $c_{9}$ is extracted by matching the diffusion-model reconstruction of $A(t)$ to the $\mathcal{O}(p^{4})$ ChPT expression $A(t)=a(t)+c_{9}\,b(t)$ via generalised least squares (GLS) at five points in $t\in[0.02,\,0.10]\;\text{GeV}^{2}$, where the chiral truncation error remains below $1.3\%$. Using the diffusion-model covariance as the GLS weight matrix, we obtain
$c_{9}=-0.61\,^{+0.17}_{-0.17}\;\text{(stat.)}\;\pm\,0.05\;\text{(ChPT trunc.)}\;\pm\,0.06\;\text{(fit range)}\,.$ (15)
@@
A cross-check using $J(t)$ at $\mathcal{O}(p^{3})$ yields $c_{9}\big|_{J}=-0.40\,^{+0.41}_{-0.41}$, consistent within uncertainties. For the physical-pion-mass analysis we combine the uncertainties in quadrature and adopt $c_{9}=-0.61\pm 0.19$. Similarly, $c_{8}$ is extracted from $D(t)$ at $\mathcal{O}(p^{2})$ via $D(t)=a(t)+c_{8}\,b(t)$ at four points in $t\in[0.02,\,0.08]\;\text{GeV}^{2}$:
$c_{8}=-4.6\,^{+0.6}_{-0.6}\;\text{(stat.)}\;\pm\,0.3\;\text{(ChPT trunc.)}\;\pm\,0.4\;\text{(fit range)}\,.$ (16)
@@
This yields $c_{8}=-4.6\pm 0.8$ (all uncertainties added in quadrature). The larger reconstruction uncertainty near $t=0$ ($\sigma_{\mathrm{diff}}\sim 0.8$$1.4$) reflects the absence of a sum-rule constraint on $D(0)$ and the large errors existing in the Lattice QCD at the unphysical pion mass around $t\approx 0$. A sharper determination of $D(t)$ with the present lattice data would require including the full set of generally covariant chiral Lagrangian contributions up to $\mathcal{O}(p^{4})$, incorporating explicit $\Delta$-resonance degrees of freedom, and performing a dedicated analysis to determine the additional low-energy constants that enter at this order Alharazin:2023uhr .
@@
It is instructive to compare these values with the model-independent determination by Cao et al. Cao:2025dkv , who employ a data-driven dispersive framework in which the nucleon GFF spectral functions are constructed from $\pi\pi$ and $K\bar{K}$ intermediate states using Roy-Steiner equation solutions for the $\pi\pi/K\bar{K}\to N\bar{N}$ partial-wave amplitudes, matched to the ChPT expressions with external gravitational source. They obtain $c_{8}=-4.28\,^{+0.37}_{-0.38}\penalty 10000\ \mathrm{GeV}^{-1}$ and $c_{9}=-0.68\,^{+0.06}_{-0.05}\penalty 10000\ \mathrm{GeV}^{-1}$ Cao:2025dkv . Both values are in excellent agreement with our extractions, $c_{8}=-4.6\pm 0.8$ and $c_{9}=-0.61\pm 0.19$, despite the two approaches being methodologically entirely distinct: theirs reconstructs spectral functions from scattering data at the physical pion mass, while ours conditions a diffusion-model posterior on lattice QCD results at an unphysical pion mass. That the central values agree well within one standard deviation for both LECs provides a non-trivial cross-check for both determinations. In Appendix D, the extracted values of $c_{8}$ and $c_{9}$ are substituted into the ChPT expressions for the nucleon GFFs, which are then evaluated at the physical pion mass to provide approximate estimates, with uncertainties reflecting the limited high-$|t|$ information.
V Summary and Outlook

### v2

@@
We develop a generative framework based on denoising diffusion for the model-independent reconstruction of hadronic form factors from sparse and noisy data. The generative prior is built from a large ensemble of synthetic curves drawn from ten distinct functional classes rooted in different theoretical approaches to hadron structure. Applied to the proton gravitational form factors $A(t)$, $J(t)$, and $D(t)$, the framework yields non-parametric reconstructions consistent with lattice QCD across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$, remaining robust even when only one or two conditioning points are retained. The densely sampled output enables a direct extraction of the chiral low-energy constants $c_{8}=-4.6\pm 0.8\ \rm GeV^{-1}$ and $c_{9}=-0.61\pm 0.19\ \rm GeV^{-1}$. Using these values at the physical pion mass, we obtain $D(0)=-4.3\pm 0.8$ for the nucleon $D$-term.
@@ I Introduction
The hadronic matrix elements of the energy-momentum tensor (EMT) have been studied intensively over the past decades, both theoretically and experimentally. While much attention has been devoted to the physical interpretation of the Fourier transforms of the gravitational form factors (GFFs) in terms of spatial density distributions of energy (mass), spin, pressure, and shear forces inside hadrons, see, e.g., Jaffe:2020ebz ; Panteleeva:2022uii ; Freese:2021mzg ; Ji:2025qax , the GFFs themselves, as functions of momentum transfer, are interesting in their own right. In particular, they can be extracted from generalized parton distributions (GPDs) Muller:1994ses ; Ji:1996ek ; Radyushkin:1996nd ; Goeke:2001tz , which are experimentally accessible through exclusive processes such as deeply virtual Compton scattering (DVCS) Ji:1996ek ; Radyushkin:1997ki ; Burkert:2018bqq ; Burkert:2021ith and hard exclusive meson production Collins:1996fb . Knowing the GFFs of a system that can be probed experimentally, such as the nucleon, establishes a direct connection between theoretical predictions and experiment, thereby testing the validity and consistency of the underlying theory.
@@
The remainder of this paper is organised as follows. In Sec. II we describe the construction of the training dataset from ten classes of physically motivated functional forms and their convex combinations. Section III presents the DDPM architecture, including the $v$-prediction formulation, the dual conditioning mechanism, and the error-propagation scheme. In Sec. IV we demonstrate the model-independent reconstruction of the proton GFFs at unphysical pion mass and extract the LECs $c_{8}$ and $c_{9}$. In Sec. V we reconstruct the GFFs at physical pion mass. Network architecture, training procedure, ablation studies on synthetic test curves are collected in Appendices A and B.
@@ III.3 Conditioning and error propagation
Our goal is to reconstruct the full GFF curve given $K$ known data points [^3] Here $F_{k}$ refers to a (gravitational) form factor $F(t_{k})$. $\{(t_{k},\,F_{k}\pm\sigma_{k})\}_{k=1}^{K}$ from, e.g., lattice QCD or ChPT. The model conditions on this information through two complementary mechanisms. First, we have the concatenation conditioning: each known value is mapped to the nearest grid index $j_{k}$ and normalised to the model’s internal representation: $\tilde{F}_{k}=(F_{k}-\mu_{j_{k}})/\sigma_{j_{k}}$, where $\mu$ and $\sigma$ are the per-grid-point mean and standard deviation of the training set. The network receives three input channels at every reverse step: the current noisy estimate $x_{\tau}$, a binary mask $m\in\{0,1\}^{L}$ with $m_{j_{k}}=1$ at known positions, and a condition vector $c\in\mathbb{R}^{L}$ with $c_{j_{k}}=\tilde{F}_{k}$ and zero elsewhere. During training, random masks (50% uniformly random, 30% clustered in the low-$|t|$ region, 20% unconditional) teach the network to reconstruct missing values conditioned on an arbitrary subset of known points. To propagate the uncertainties of the known data into the reconstructed uncertainty band, each of the $N$ generated samples receives its own realisation of the conditioning values [^4] It is worth mentioning that throughout the lattice reconstructions presented in the next section, the uncertainties of the conditioning points are propagated independently via Eq. (11), as the full covariance matrix of Ref. Hackett:2023nkr is publicly not available; the impact of off-diagonal elements is expected to be subdominant, since the reconstruction remains stable under data reduction, as we will see below.,
@@ IV Model-independent reconstruction of the proton GFFs at unphysical pion mass
In this section, we demonstrate that the diffusion-based generative framework yields a fully model-independent reconstruction of the total proton GFFs $A(t)$, $J(t)$, and $D(t)$, consistent with the lattice QCD results of Ref. Hackett:2023nkr across the full kinematic range $0\leq-t\leq 2\penalty 10000\ \mathrm{GeV}^{2}$, without imposing any functional form on the $t$-dependence. The only inputs to the reconstruction are:
@@
Having learned the space of physically admissible form-factor shapes from the training prior (Sec. II), the diffusion model samples from the posterior distribution consistent with the supplied conditioning input, yielding a non-parametric estimate of both the central value and the pointwise uncertainty of the reconstructed form factor.
@@
We emphasize that for each available lattice dataset we select as conditioning input those points with only modest statistical uncertainties compared with the others, in other words, the reconstruction is not aided by large error bars that would trivially accommodate many functional shapes. The diffusion model must reproduce sharply defined values around the conditioning points while simultaneously generating physically sensible interpolations and extrapolations throughout the kinematic range. That it succeeds with as few as one or two data points is a direct consequence of the physically informed prior built into the training data, and constitutes a stringent validation of the approach.
@@ IV.1 Reconstruction of symmetry-constrained form factors $A(t)$ and $J(t)$
  • Reconstruction from minimal input: The most notable result appears in the bottom panels, where only a single lattice data point at intermediate t-t and the forward-limit constraint are supplied. Even in this extreme scenario the reconstruction yields a smoothly decaying curve whose median remains compatible with the full parametric fits. At $t=2\penalty 10000\ \mathrm{GeV}^{2}$ the credible band remains moderate for both form factors. The notably narrow uncertainty obtained with only two conditioning inputs, despite the wide range of mathematically allowed curves through those points, is a direct consequence of the physical prior encoded in the training data.
@@ IV.2 Reconstruction of D form factor
  • Minimal constraint ($D(0)<0$): The requirement that $D(0)$ be negative follows from the mechanical stability of the proton Polyakov:2018zvc ; Perevalova:2016dln : a positive $D$-term would imply a repulsive normal force and the system would collapse. This negativity is also supported by ChPT Gegelia:2021wnj , chiral quark-soliton models Goeke:2007fp ; Cebulla:2007ei , and the lattice data themselves. This weakest condition allows the form factor to fluctuate mildly at large t-t.
  • Full negativity ($D(t)<0$): Negativity is enforced at every point on the spacelike grid. This stronger condition is motivated by all known model calculations, see e.g. Refs. Cebulla:2007ei ; Cao:2025ncp ; Yao:2024ixu ; Cao:2025dkv ; Deng:2025azp .
@@
As a further application of our approach, we reconstruct the quark and gluon contributions to the $D(t)$ form factor using two lattice data points from Ref. Hackett:2023nkr and compare our result with the model prediction of Ref. Guo:2025jiz . As can be seen from Figs. (3(a)) and (3(b)), the model-independent reconstruction from only two lattice data points is consistent with the tripole parameterization of Ref. Guo:2025jiz . This is again noteworthy because our method does not assume a specific functional form of the GFF.
Abbildung 3: Diffusion-model reconstruction of $D(t)$ at the unphysical pion mass, conditioned on the lattice data from Ref. Hackett:2023nkr , compared with the model prediction from Ref. Guo:2025jiz .
@@ IV.3 Extracting the unknown LECs $c_{8}$ and $c_{9}$
The low-energy constant $c_{9}$ is extracted by matching the diffusion-model reconstruction of $A(t)$ to our $\mathcal{O}(p^{4})$ ChPT expression $A(t)=a(t)+c_{9}\,b(t)$ as derived in Refs. Alharazin:2020yjv ; Alharazin:2023uhr , using generalised least squares (GLS) at five points in $t\in[0.02,\,0.10]\;\text{GeV}^{2}$, where the chiral truncation error remains below $1.3\%$. Using the diffusion-model covariance as the GLS weight matrix, we obtain
$c_{9}=-0.61\,^{+0.17}_{-0.17}\;\text{(stat.)}\;\pm\,0.05\;\text{(ChPT trunc.)}\;\pm\,0.06\;\text{(fit range)}\ \rm GeV^{-1}\,.$ (15)
@@
A cross-check using $J(t)$ at $\mathcal{O}(p^{3})$ yields $c_{9}\big|_{J}=-0.40\,^{+0.41}_{-0.41}\ \rm GeV^{-1}$, consistent within uncertainties. For the physical-pion-mass analysis we combine the uncertainties in quadrature and adopt $c_{9}=-0.61\pm 0.19\ \rm GeV^{-1}$. Similarly, $c_{8}$ is extracted from $D(t)$ at $\mathcal{O}(p^{2})$ via $D(t)=a(t)+c_{8}\,b(t)$ at four points in $t\in[0.02,\,0.08]\;\text{GeV}^{2}$:
$c_{8}=-4.6\,^{+0.6}_{-0.6}\;\text{(stat.)}\;\pm\,0.3\;\text{(ChPT trunc.)}\;\pm\,0.4\;\text{(fit range)}\ \rm GeV^{-1}\,.$ (16)
@@
This yields $c_{8}=-4.6\pm 0.8\ \rm GeV^{-1}$ (all uncertainties added in quadrature). The larger reconstruction uncertainty near $t=0$ ($\sigma_{\mathrm{diff}}\sim 0.8$$1.4$) reflects the absence of a sum-rule constraint on $D(0)$ and the large errors existing in the Lattice QCD at the unphysical pion mass around $t\approx 0$. A sharper determination of $D(t)$ with the present lattice data would require including the full set of generally covariant chiral Lagrangian contributions up to $\mathcal{O}(p^{4})$, incorporating explicit $\Delta$-resonance degrees of freedom, and performing a dedicated analysis to determine the additional low-energy constants that enter at this order Alharazin:2023uhr .
@@
It is instructive to compare these values with the model-independent determination by Cao et al. Cao:2025dkv , who employ a data-driven dispersive framework in which the nucleon GFF spectral functions are constructed from $\pi\pi$ and $K\bar{K}$ intermediate states using Roy-Steiner equation solutions for the $\pi\pi/K\bar{K}\to N\bar{N}$ partial-wave amplitudes, matched to the ChPT expressions with external gravitational source. They obtain $c_{8}=-4.28\,^{+0.37}_{-0.38}\penalty 10000\ \mathrm{GeV}^{-1}$ and $c_{9}=-0.68\,^{+0.06}_{-0.05}\penalty 10000\ \mathrm{GeV}^{-1}$ Cao:2025dkv . Both values are in excellent agreement with our extractions, $c_{8}=-4.6\pm 0.8\ \rm GeV^{-1}$ and $c_{9}=-0.61\pm 0.19\ \rm GeV^{-1}$, despite the two approaches being methodologically entirely distinct: theirs reconstructs spectral functions from scattering data at the physical pion mass, while ours conditions a diffusion-model posterior on lattice QCD results at an unphysical pion mass. That the central values agree well within one standard deviation for both LECs provides a non-trivial cross-check for both determinations. In the next section, the extracted values of $c_{8}$ and $c_{9}$ are substituted into the ChPT expressions for the nucleon GFFs, which are then evaluated at the physical pion mass to provide approximate estimates, with uncertainties reflecting the limited high-$|t|$ information.
V The proton GFFs at physical pion mass
V.1 Reconstruction of GFFs from ChPT
Figures (4) and (5) presents the reconstruction of all three GFFs at the physical pion mass. For each form factor, the three panels show the effect of progressively reducing the number of ChPT conditioning points, from the full set (top) to a minimal subset (bottom). For $A(t)$ and $J(t)$ (Figs. (4(a)) and (4(b))), the Poincaré constraints at $t=0$ anchor the normalisation, and the reconstruction remains stable even as conditioning points are removed.
The form factor $D(t)$ (Fig. (5(a))) presents a qualitatively different situation. Unlike $A(t)$ and $J(t)$, no sum-rule constraint fixes $D(0)$, so the reconstruction relies entirely on the ChPT conditioning data. Moreover, the available ChPT expression for $D(t)$ is of lower chiral order than those for $A(t)$ and $J(t)$, so that the higher-$|t|$ conditioning points probe a regime where the truncated chiral expansion is no longer reliable and the quoted uncertainties underestimate the true error. The data-ablation sequence reveals that removing these higher-$|t|$ points actually improves the reconstruction: the diffusion-model median is not pulled toward the potentially unreliable high-$|t|$ ChPT values but instead follows a trajectory shaped by the training prior, effectively filtering out the incomplete chiral information.
$D(0)\;=\;c_{8}\,m_{N}\;=\;-4.3\;\pm\;0.8\,$, (17)
It is worth noting that for the form factor $D(t)$ plotted in Fig. (5(a)), we employed the unexpanded ChPT results of Refs. Alharazin:2020yjv ; Alharazin:2023uhr . However, these results do not yet include all contributions to $D(t)$ arising from the third- and fourth-order chiral Lagrangians. In particular, the fourth-order Lagrangian in covariant form has yet to be constructed in its most minimal and general form. For the determination of the $D$-term, we therefore restricted ourselves to the low-energy region and retained only the lowest-order terms from the small-quantity expansion of the results in Refs. Alharazin:2020yjv ; Alharazin:2023uhr . This also explains the difference between Fig. (6), where we used the complete expanded ChPT results at lowest order, and Fig. (5(a)), where we used the unexpanded but incomplete ChPT contributions.
Abbildung 4: Diffusion-model reconstruction of $A(t)$ and $J(t)$ at the physical pion mass, conditioned on ChPT data points. Panel layout follows the data-ablation protocol, with fewer conditioning points from top to bottom
Abbildung 5: Diffusion-model reconstruction of $D(t)$ at the physical pion mass, conditioned on ChPT data points. Panel layout follows the data-ablation protocol, with fewer conditioning points from top to bottom
V.2 Reconstruction of GFFs from DVCS data
In Ref. Burkert:2021ith , the $D(t)$ form factor was extracted from deeply virtual Compton scattering data measured with the CLAS detector at Jefferson Lab. It should be noted, however, that this extraction relies on specific parameterizations of the Compton Form Factors and the assumption of GPD H dominance, rendering the resulting $D(t)$ values inherently model dependent Kumericki:2019ddg ; Dutrieux:2021nlz . Nevertheless, these data points constitute the only currently available experimental determination of the gravitational form factor $D(t)$ of the proton. In the following, we reconstruct the $D(t)$ form factor from the extracted values reported in Ref. Burkert:2021ith in two scenarios: first, using the data points alone, and second, incorporating the forward-limit constraint obtained in the previous section. The results of the two reconstruction scenarios are shown in Fig. (6). When only the DVCS data points are used the extrapolation toward the forward limit is largely unconstrained, as reflected by the rapidly widening 68% confidence interval and the significant deviation between mean and median at low t-t. Incorporating the leading-order ChPT constraint dramatically reduces the uncertainty in the forward region.
Abbildung 6: Diffusion-model reconstruction of $D(t)$ at the physical pion mass, conditioned only on the DVCS data from Burkert:2021ith and combined with the ChPT forward limit obtained in Sec. V.1.
VI Summary and Outlook